Please help to show the work.

(1)Compute (a) (f+g)x, (b)(f*g)x, and f(g(x)) given f(x)= x^2-1 and g(x)=x+2.

Answer by Nate(3500)

f(x) = x^2 - 1 and g(x) = x + 2
(f + g)x = x^2 - 1 + x + 2
(f + g)x = x^2 + x + 1
(f * g)x = (x^2 - 1)(x + 2)
(f * g)x = x^3 + 2x^2 - x - 2
f[g(x)] = (x + 2)^2 - 1
f[g(x)] = x^2 + 4x + 3